OpenStudy (anonymous):

The volume of a sphere is 4,000 m3. What is the surface area of the sphere to the nearest square meter @Directrix

Directrix (directrix):

We have both volume and surface area this time.

Directrix (directrix):

The only thing to know about a sphere is its radius. So, we will have to get that first.

Directrix (directrix):

V = 4/3 * π * r³ 4,000 = 4/3 * 3.14 * r³ 3000/3.14 = r³ Solve for r

OpenStudy (anonymous):

kinda confused by all the numbers

Directrix (directrix):

Yes, I used the online calculator You can see the result at this link: http://www.wolframalpha.com/input/?i=4%2C000+%3D+4%2F3+*+pi+*+r%C2%B3

Directrix (directrix):

r = 9.84

Directrix (directrix):

Now, to get surface area. A = 4 * π * r² A = 4 * π * (9,84)²

OpenStudy (anonymous):

approximate the pi

Directrix (directrix):

Yes. Approximate the pi unless the instructions tell you to leave the answer in terms of pi.

Directrix (directrix):

1216.75 which to the nearest square meter is 1217 square meters. Looks a bit odd.

OpenStudy (anonymous):

no it doesn't say that so i guess i do

OpenStudy (anonymous):

the closest to it is 1219

Directrix (directrix):

We should not be that far off. Let me see.

Directrix (directrix):

I allowed for more digits in the approximation of r. A= 4 * π * (9.8475)² A = 1218.6 so that goes to 1219

OpenStudy (anonymous):

i got it ?

OpenStudy (anonymous):

1219m^2

Directrix (directrix):

When you are taking the timed test, if you get close to the correct answer, don't go back to add extra digits of a number. Because on multiple choice, go with anything close.

OpenStudy (anonymous):

good advice ,ill keep that in mind

OpenStudy (anonymous):

ready for the last one?

Directrix (directrix):

Yes.